OpenStudy (anonymous):

divide f(x) by d(x) and write a summary statement in the form indicated f(x)=x^4+5x^3-2x^2+5x-3 ; d(x)=x^2+1

OpenStudy (anonymous):

f(x)/d(x) = (x^4 + 5x^3 - 2x^2 + 5x-3) / (x^2 + 1) you have to do this by long division: .........................x^2 + 5x - 3 ............ __________________________ x^2 + 1 ) x^4 + 5x^3 - 2x^2 + 5x - 3 ............-(x^4 + 0x^3 + x^2) <---Note this is (x^2 +1)*x^2 ......................5x^3 - 3x^2 + 5x ....................-(5x^3 +0x^2 + 5x) <---this is (x^2 +1)*5x ...............................-3x^2 + 0x - 3 .............................-(-3x^2 + 0x - 3) <and this is (x^2 +1)*(-3) ......................................... So after dividing, you end up with: x^2 + 5x -3

OpenStudy (anonymous):

@psila hope this helps...

OpenStudy (anonymous):

hello? @psila

OpenStudy (anonymous):

ah! forget it!!

OpenStudy (anonymous):

sorry i don't understand

OpenStudy (anonymous):

it's also not one of the answers

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@raster

OpenStudy (anonymous):

are there some choices?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

f(x)=(x^2+1)(x^2+5x-3) f(x)=(x^2+1)(x^2+5x-3)-15x-9 f(x)=(x^2+1)(x^2-5x-3)-15x-9 f(x)=(x^2+1)(x^2-5x-3)

OpenStudy (anonymous):

a and d are the same.....and so are b and c

OpenStudy (anonymous):

is this from an online school?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it's a after i distributed

OpenStudy (anonymous):

i got the original problem. they have different signs

OpenStudy (anonymous):

ok....follow my steps..... 1. take a screenshot of the question by pressing print scrn on your keyboard

OpenStudy (anonymous):

2. open paint and press ctrl + v at the same time...

OpenStudy (anonymous):

3. go back to openstudy and reply...

OpenStudy (anonymous):

thanks! i will try it!

OpenStudy (anonymous):

4. click on attach file...make sure you have the screenshot saved

OpenStudy (anonymous):

there you go.....

OpenStudy (anonymous):

@psila got it?

OpenStudy (anonymous):

@raster